A 22.5 g sample of ammonium carbonate contains___________ mol of ammonium ions.
Answer #1

Ammonium carbonate is (NH4)2CO3, this will produce 2 moles of NH4 (ammonium ions)

calculate the moles ammonium carbonate

moles = mass / molar mass, molar mass of ammonium carbonate = 96.09

moles = 22.5 / 96.09 = 0.234 moles

so you will have 2 * 0.234 = 0.468 moles of ammonium ions

*hope this helps =)